Two spherical objects have masses of 66g and 35 g. The gravitational attraction between them is 6.5 x 10-3 N. How far apart are their centers?


1
Expert's answer
2022-04-06T13:57:09-0400

Given:

"m_1=0.066\:\rm kg"

"m_2=0.035\:\rm kg"

"F=6.5*10^{-3}\:\rm N"


The Newton's law of universal gravitation says

"F=G\frac{m_1m_2}{r^2}"

Hence, the distance between bodies

"r=\sqrt{G\frac{m_1m_2}{F}}"

"r=\sqrt{6.67*10^{-11}\frac{0.066*0.035}{6.5*10^{-3}}}=4.9*10^{-6}\rm\: m"